a lottery consists of 10000 tickets out of which 10 are prizes. if a person selects 2 tickets what is the probability of not getting a prize
1
Expert's answer
2013-02-05T10:01:57-0500
The probability that the ticket& is prize is

P(the ticket is prize) = 10/10000 = 0.0001.

If a person has two tickets then the probability to win the prize is

P(to win the prize with two tickets) = 2·P(the ticket is prize) = 0.0002.

Therefore, the probability of not getting a prize is

P(the probability of not getting a prize with two tickets) =

1 - P(to win the prize with two tickets) =

1 - 0.0002 = 0.9998.